Please prove the following statement if it is true, or provide counter example if it is false:

In a linear model, the ordinary least squares residual vector હૈ := (€ ... 6) satisfies that XT? for all j = 1,...,p, where X_j is the jth column of X
Answer #1

we krea, X+ Model -1 tne LSEAadorfr is f-(XTx) y pTNided (7x)i ther yxp X (NXTY 13 2 = Y- Y - x(xxxTy - (1- xxY (21 2 ) =jthTo anoeo 2 t is t 8how that =0 5-101p Cy-25 0 5= 1CUP 5-IC1OP Y(1-xixixixT) : I-Xi)XT1 X X-X(XTxTx) = x-X O y (1-X(XTXT) (yPlease LIKE!!!